### v4

# Learning Mamba as a Continual Learner: Meta-learning State Space Models for Efficient Continual Learning

Abstract
Continual learning (CL) aims to efficiently learn from a non-stationary data stream, without storing or recomputing all seen samples. CL enables prediction on new tasks by incorporating sequential training samples. Building on this connection between CL and sequential modeling, meta-continual learning (MCL) aims to meta-learn an efficient continual learner as a sequence prediction model, with advanced sequence models like Transformers being natural choices. However, despite decent performance, Transformers rely on a linearly growing cache to store all past representations, conflicting with CL’s objective of not storing all seen samples and limiting efficiency. In this paper, we focus on meta-learning sequence-prediction-based continual learners without retaining all past representations. While attention-free models with fixed-size hidden states (e.g., Linear Transformers) align with CL’s essential goal and efficiency needs, they have shown limited effectiveness in MCL in previous literature. Given Mamba’s strong sequence modeling performance and attention-free nature, we explore a key question: Can attention-free models like Mamba perform well on MCL? By formulating Mamba and the SSM for MCL tasks, we propose MambaCL, a meta-learned continual learner. To enhance MambaCL’s training, we introduce selectivity regularization, leveraging the connection between Mamba and Transformers to guide its behavior over sequences. Furthermore, we study how Mamba and other models perform across various MCL scenarios through extensive and well-designed experiments. Our results highlight the promising performance and strong generalization of Mamba and attention-free models in MCL, demonstrating its potential for efficient continual learning and adaptation.
1 Introduction
Continual learning (CL) aims to efficiently learn and accumulate knowledge in a non-stationary data stream containing different tasks, without storing or recomputing training on all previously seen data [11, 64]. Given a sequence of data ${\mathcal{D}}_{T}=(({\mathbf{x}}_{1},y_{1}),...,({\mathbf{x}}_{t},y_{t}),...,({\mathbf{x}}_{T},y_{T}))$ with a series of paired observations ${\mathbf{x}}_{i}$ (e.g., images) and targets $y_{i}$ (e.g., class labels) from different tasks, CL is usually produced to learn one model $P_{\phi_{t}}(y|{\mathbf{x}})$ parameterized by $\phi_{t}$ that can perform prediction for any tasks corresponding to the seen data ${\mathcal{D}}_{t}$. For example, in class incremental learning (CIL) [49, 73], a widely studied CL scenario, ${\mathcal{D}}_{T}$ consists of data with incrementally added classes, and $P_{\phi_{t}}(y|{\mathbf{x}})$ is trained to recognize all seen classes. To ensure efficiency, CL methods learn from data streams while minimizing historical data storage and constraining memory growth, ideally at least a sub-linear rate [11, 44, 63]. The key challenge in CL is maintaining performance on past tasks while continuously updating model parameters $\phi_{t}$ [11, 64].
CL methods continually train/update the model $P_{\phi_{t}}(y|{\mathbf{x}})$ from seen sequence ${\mathcal{D}}_{t}$ at arbitrary step $t$, enabling prediction of the output $y^{\text{test}}$ for any observation ${\mathbf{x}}^{\text{test}}$. The whole learning and prediction process can be viewed as a sequence prediction (SP) problem [32, 3], where the goal is to predict $y_{\text{test}}$ for a query ${\mathbf{x}}_{\text{test}}$ conditioning on the seen data sequence and the testing input, i.e., $({\mathcal{D}}_{t}^{\text{train}},{\mathbf{x}}^{\text{test}})\equiv({\mathbf{x}}_{1}^{\text{train}},y_{1}^{\text{train}},...,{\mathbf{x}}_{t}^{\text{train}},y_{t}^{\text{train}},{\mathbf{x}}^{\text{test}})$. In conventional CL, the model parameter $\phi_{t}$ is trained to maintain the historical knowledge (as a states on the sequence) through an optimization step $\phi_{t+1}=\texttt{optim-step}(\phi_{t},{\mathbf{x}}_{t},y_{t})$. This connection between SP and CL motivates meta-learning a continual learner as a sequence prediction model, for computation-and-data-efficient CL. Through meta-continual learning (MCL) framework [32, 58, 3], a continual learner $f_{\theta}()$ parameterized by $\theta$ can be trained multiple CL episodes. A meta-learned $f_{\theta}()$ can take a sequence $({\mathcal{D}}_{t},{\mathbf{x}}_{t+1})$ as input and predict $y_{t+1}=f_{\theta}(({\mathcal{D}}_{t},{\mathbf{x}}_{t+1}))$, equivalent to a predictive model $P_{\theta}(y|{\mathbf{x}},{\mathcal{D}}_{t})$. The data stream serves as a context for making predictions on new queries.
Considering the strong sequence modeling ability, Transformers [65, 61] are a natural choice for SP-based MCL [32, 58]. In MCL, given a data stream, a meta-learned Transformer generates a new key-value pair at each step and maintains the key-value pairs for all seen training samples in a key-value cache. The prediction for the query at each step is based on attention over the key-value pairs retained from all preceding training samples. This retrieval-based approach enables strong performance in CL and MCL datasets [32, 3]. Benefiting from the retrieval-based modeling and explicit intra-sequence attention, Transformers can perform effectively on the datasets formulated for CL and MCL [32, 3]. However, Transformers rely on maintaining the key-value cache of all seen samples. While the key-value cache resembles a recurrent neural network (RNN) hidden state [27, 32], its size grows linearly with seen tokens, increasing memory and computation costs, which contradicts CL’s goal of learning without storing full historical data. The basic Transformer may not fully fit the intention of MCL, despite the reported decent performance [32].
In this paper, we focus on MCL without retaining all past representations. Among various approaches to improving Transformer efficiency [60], attention-free models achieve efficiency by approximating softmax attention with kernel methods and linear operations, resulting in constant hidden state sizes and linear computation complexity, e.g., Linear Transformer (Linear TF) [27] and Performer [8]. While these models align better with CL’s goals, they struggle in MCL due to limited expressive power [32, 27, 8, 60]. Recent advances in state-space models (SSMs) have led to efficient attention-free models for long-sequence processing with nearly linear computation [17, 18]. By incorporating time-varying modeling, Mamba [16, 9] achieves outstanding performance in language modeling [15].
Given Mamba’s strong performance as an attention-free model, we ask: Can the attention-free model like Mamba perform well in MCL? To explore this, we propose MambaCL, formulating selective SSM for MCL and meta-learning Mamba as an online continual learner. Training SP models for MCL proves challenging, especially for Mamba, which lacks explicit attention operations as in Transformers. To address this, we introduce a selectivity regularizer leveraging connections between SSM/Mamba, Linear Transformers, and Transformers for training, guiding the learning of time-variant operations and behaviors in the selective SSM during training. Beyond the scope of the existing works [32] focusing on basic MCL formulation and setting, we expand the studies to more realistic scenarios and try to answer: how different models (including Transformers and Mamba) perform across various MCL scenarios. Our experiments show that Mamba can perform well in most MCL scenarios, significantly outperforming other attention-free methods like Linear Transformers and matching or surpassing Transformers with fewer parameters and computations. Notably, in some challenging settings with strong global structures (e.g., fine-grained) and scenarios with domain shifts or long sequences, Mamba demonstrates promising reliability, generalization, and robustness, highlighting its potential for efficient CL and adaptation. We also analyze model design influences and explore potential Mamba variants, such as incorporating mixture-of-experts for MCL.
2 Related Work
Continual learning aims to address catastrophic forgetting when training models across sequential tasks [11, 64]. Existing approaches include replay, regularization, and architecture-based methods. Replay-based techniques retain task knowledge via memory buffers [49, 37, 7, 6] or generate pseudo-samples using generative models [57, 52, 51], though they face memory and privacy constraints. Regularization-based methods preserve previous knowledge by penalizing important parameter changes or applying distillation [28, 70, 43, 34, 1, 71]. Architecture-based methods mitigate interference by assigning task-specific parameters via masking or dynamic adaptation [69, 56, 33, 67, 68, 63, 38].
Meta-learning aims to learn generalizable models across tasks, enabling efficient adaptation to new tasks with limited data. The meta-learning is inherently suited for and has been adapted for CL, focusing on balancing the fit for current tasks while maintaining generalization across all previously encountered tasks [50, 2, 21, 66, 41, 42, 55, 20, 4, 39, 23]. Meta-continual learning (MCL) deviates from traditional CL settings by incorporating multiple CL episodes, structured into meta-training and meta-testing sets [58, 32, 3]. [32] frames MCL as a sequence modeling task, aligning continual learning with language modeling. OML [25] adopts a dual-architecture, updating predictor while freezing encoder during training, then jointly optimizing both during meta-testing for stability.
Figure 1: The overall framework of our proposed methods. We meta-train a Mamba Learner $f_{\theta}()$ to perform meta-continual learning (MCL) by processing an online data stream containing paired $({\mathbf{x}},y)$ examples. Meta-learning of this continual learner is conducted across multiple CL episodes. The model produces predictions by relying on the retained hidden state. Here, we demonstrate how the Mamba learner recurrently processes input data at steps 00, $2$, and $t-1$, respectively.
Transformer architectures excel in sequence modeling due to their attention mechanism [65]. Decoder-only variants like GPT [5] and LLaMA [61], which process inputs causally, have driven major advances in deep learning. However, their softmax-based attention incurs quadratic complexity with sequence length, limiting scalability. This has spurred interest in RNN-like alternatives with linear-time modeling. In particular, linear attention methods replace the exponential similarity with dot products over transformed key/query vectors, offering improved efficiency [27, 8, 59].
State Space Models (SSMs), inspired by classical control theory [26], have recently gained traction for sequence modeling [17, 18]. Mamba advances this framework by introducing time-varying parameters and a hardware-efficient implementation [16, 9], achieving fast training and inference. It has proven effective in vision and language tasks [74, 72, 22, 35].
3 Problem Formulation and Methodology
In Continual Learning (CL), given a non-stationary data stream ${\mathcal{D}}_{T}^{\text{train}}=\{({\mathbf{x}}_{t},y_{t})\}_{t=1}^{T}$ with ${\mathbf{x}}_{t}\in{\mathcal{X}}_{t}$ and $y_{t}\in{\mathcal{Y}}_{t}$ as training data, a predictive model $g_{\phi_{t}}():{\mathcal{X}}\rightarrow{\mathcal{Y}}$ is trained on the stream as $P_{\phi_{t}}(y|{\mathbf{x}})$ (at step $t$) for a potential testing set ${\mathcal{D}}_{T}^{\text{test}}=\{({\mathbf{x}}_{n},y_{n})\}_{n=1}^{N}$, where ${\mathbf{x}}_{t}\in{\mathcal{X}}_{t}$ and $y_{t}\in{\mathcal{Y}}_{t}$, following the distribution of the training set. A conventional continual learner is manually crafted to continually update model parameters $\phi_{t}$ through optimization. The data stream ${\mathcal{D}}_{T}$ typically comprises samples from different tasks or distributions, often piecewise stationary within a task interval. In online CL, each sample is seen only once, whereas in offline CL, task-specific samples can be accessed for multiple rounds. This work primarily focuses on the online CL setting.
For achieving efficient CL, Meta-Continual Learning (MCL) [32, 58] meta-learns a parameterized continual learner that updates a predictive model (in ${\mathcal{X}}\rightarrow{\mathcal{Y}}$) from a streaming data source (in ${\mathcal{X}}\times{\mathcal{Y}}$). Given a learned model from a sequence ${\mathcal{D}}_{t}^{\text{train}}$ in MCL, prediction can be formulated as $P_{\theta}(y_{\text{test}}|{\mathbf{x}}_{\text{test}},{\mathcal{D}}_{t}^{\text{train}})$. MCL can be seen as training a sequence prediction (SP) model $f_{\theta}():({\mathcal{X}}\times{\mathcal{Y}})\times{\mathcal{X}}\rightarrow{\mathcal{Y}}$, where $f_{\theta}()$ continually processes streaming data and performs prediction for test samples via $\hat{y}_{\text{test}}=f_{\theta}({\mathbf{x}}_{\text{test}},{\mathcal{D}}_{t}^{\text{train}})$. The learner updates its hidden states dynamically, reflecting the continual learning process. Through multiple episodes of $({\mathcal{D}}^{\text{train}},{\mathcal{D}}^{\text{test}})$, the learner’s parameters $\theta$ are optimized via meta-learning to improve performance across all. The targets $y$ in different episodes are independent, not sharing semantic meaning. This work focuses on MCL using a parameterized SP model, despite other meta-learning approaches [14, 25].
3.1 Preliminaries: Transformers and SSMs
Transformers produce next-token predictions in sequence relying on a self-attention mechanism [65]. Given a sequence of $N$ vectors in $M$-dimension denoted as ${\mathbf{Z}}\in{\mathbb{R}}^{N\times M}$, the vanilla self-attention is formulate with a softmax attention method:
$\begin{split}{\mathbf{Q}}\!={\mathbf{Z}}{\mathbf{W}}_{Q},~{}{\mathbf{K}}\!={\mathbf{Z}}{\mathbf{W}}_{K},~{}{\mathbf{V}}\!={\mathbf{Z}}{\mathbf{W}}_{V},~{}~{}~{}{\mathbf{u}}_{t}=\sum_{j=1}^{N}\ \frac{\exp\left({\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}/{\sqrt{d}}\right)}{\sum_{j=1}^{N}\ \exp\left({\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}/{\sqrt{d}}\right)}{\mathbf{V}}_{j},\end{split}$ (1)
where ${\mathbf{W}}_{Q}\in{\mathbb{R}}^{M\times C},~{}{\mathbf{W}}_{K}\in{\mathbb{R}}^{M\times C},~{}{\mathbf{W}}_{V}\in{\mathbb{R}}^{M\times M}$ are the projection weight matrices, ${\mathbf{u}}_{t}\in{\mathbb{R}}^{M}$ denote the output embedding. ${\mathbf{Q}}_{t}$, ${\mathbf{K}}_{j}$, and ${\mathbf{V}}_{j}$ denote the indexed vectors in the corresponding matrices. The notation fonts are slightly abused for consistency with conventions. Each input token generates a key-value pair, leading to a linear increase in key-value cache size. Softmax attention measures the similarities between the query-key pairs, leading to ${\mathcal{O}}(N^{2})$ complexity.
Linear Transformer [27] reduces the complexity relying on a linear attention method. By applying a feature representation function $\phi()$ corresponding to a kernel for ${\mathbf{Q}}$ and ${\mathbf{K}}$, the linear attention method replaces the softmax attention with a linear operation as:${\mathbf{u}}_{t}=\sum_{j=1}^{N}\frac{{\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}}{\sum_{j^{\prime}=1}^{N}{\mathbf{Q}}_{t}{\mathbf{K}}_{j^{\prime}}^{\top}}{\mathbf{V}}_{j}=\frac{{\mathbf{Q}}_{t}\left(\sum_{j=1}^{N}{\mathbf{K}}_{j}^{\top}{\mathbf{V}}_{j}\right)}{{\mathbf{Q}}_{t}\left(\sum_{j=1}^{N}{\mathbf{K}}_{j}^{\top}\right)}$, where ${\mathbf{Q}}\!=\!\phi({\mathbf{Z}}{\mathbf{W}}_{Q}),~{}{\mathbf{K}}\!=\!\phi({\mathbf{Z}}{\mathbf{W}}_{K}),~{}{\mathbf{V}}\!\!={\mathbf{Z}}{\mathbf{W}}_{V}$, $\phi(\cdot)$ is set as $\phi({\mathbf{x}})=\mathrm{elu}({\mathbf{x}})+1$ in [27]. Performer employs $\phi({\mathbf{x}})=\exp\left({\mathbf{x}}{\mathbf{W}}_{p}-\|{\mathbf{x}}\|^{2}/2\right)$, with ${\mathbf{W}}_{p}$ comprising orthogonal random vectors [8]. The computational complexity is reduced to ${\mathcal{O}}(N)$.
In practice, the attenstion operations can be implemented in autoregressive models, where calculation of ${\mathbf{u}}_{t}$ can only see the proceeding tokens with $j\leq t$. Specifically, with the causal masking, the linear attention can be rewritten as:
$\begin{split}{\mathbf{u}}_{t}=\frac{{\mathbf{Q}}_{t}\left(\sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}{\mathbf{V}}_{j}}\right)}{{\mathbf{Q}}_{t}\left(\ \sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}}\right)}=\frac{{\mathbf{Q}}_{t}{\mathbf{S}}_{t}}{{\mathbf{Q}}_{t}{\mathbf{G}}_{t}},~{}~{}~{}{\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{{\mathbf{K}}_{t}^{\top}{\mathbf{V}}_{t}},~{}{\mathbf{G}}_{t}={\mathbf{G}}_{t-1}+{{\mathbf{K}}_{t}^{\top}},\end{split}$ (2)
where ${\mathbf{S}}_{t}=\sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}{\mathbf{V}}_{j}}$ and ${\mathbf{G}}_{t}=\sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}}$. This enables recurrent computation of causal linear attention by cumulatively updating ${\mathbf{S}}_{t}$ and ${\mathbf{G}}_{t}$, which serve as internal hidden states.
State space sequence models (SSMs) [16, 17, 9] are sequence models describing a system that maps input $z_{t}\in{\mathbb{R}}$ to output $u_{t}\in{\mathbb{R}}$ through a hidden state ${\mathbf{h}}_{t}\in{\mathbb{R}}^{C\times 1}$ in a discrete sequence applied with neural networks. Specifically, SSMs can be formulated with parameters ${\mathbf{A}}\in{\mathbb{R}}^{C\times C}$, ${\mathbf{B}}\in{\mathbb{R}}^{C\times 1}$, ${\mathbf{C}}\in{\mathbb{R}}^{1\times C}$, and $D\in{\mathbb{R}}$, as ${\mathbf{h}}_{t}={\mathbf{A}}{\mathbf{h}}_{t-1}+{\mathbf{B}}z_{t},\quad u_{t}={\mathbf{C}}{\mathbf{h}}_{t}+Dz_{t}$. We can directly formulate a discrete SSM, where the ${\mathbf{A}}$ and ${\mathbf{B}}$ are transformed from a continuous version ${\mathbf{A}}^{\prime}$ and ${\mathbf{B}}^{\prime}$ relying on a timescale parameter $\Delta\in{\mathbb{R}}$, via ${\mathbf{A}}=\exp(\Delta{\mathbf{A}}^{\prime})$ and ${\mathbf{B}}={\mathbf{A}}^{-1}({{\mathbf{A}}}-{\mathbf{I}})\cdot\Delta{\mathbf{B}}^{\prime}$. ${\mathbf{A}}$ and ${\mathbf{B}}$ perform the selection or gating in hidden state updating.
3.2 SSM and Mamba for Meta-Continual Learning
Selective SSM & Mamba in MCL. The dynamics of the basic SSM are time-invariant, limiting their ability to model complex sequences. Mamba [16] incorporates a selective SSM into the model by generating input-dependent SSM parameters to reflect the input/step-sensitive selection process. The selective SSM can be written as:
${\mathbf{h}}_{t}={\mathbf{A}}_{t}{\mathbf{h}}_{t-1}+{\mathbf{B}}_{t}z_{t},~{}~{}u_{t}={\mathbf{C}}_{t}{\mathbf{h}}_{t}+Dz_{t},$ (3)
where ${\mathbf{A}}_{t}$, ${\mathbf{B}}_{t}$, and ${\mathbf{C}}_{t}$ are produced in Mamba relying on the input token at step $t$. Unlike Transformers storing key-value pairs for all input tokens (causing linear state growth), Mamba compresses context into a fixed-size hidden state, aligning with CL’s efficiency goals.
In our MCL tasks and other practical scenarios, we need Mamba to handle the input sequence ${\mathbf{Z}}\in{\mathbb{R}}^{N\times M}$ with each token as a vector ${\mathbf{z}}_{t}\in{\mathbb{R}}^{M}$. Mamba applies the selective SSM to each dimension/channel independently with:
${\mathbf{H}}_{t}=[{\mathbf{A}}_{t,i}{\mathbf{h}}_{t-1,i}+{\mathbf{B}}_{t,i}{\mathbf{z}}_{t,i}]_{i=1}^{M},~{}{\mathbf{u}}_{t}={\mathbf{C}}_{t}{\mathbf{H}}_{t}+{\mathbf{D}}\odot{\mathbf{z}}_{t},$ (4)
where ${\mathbf{H}}_{t}\in{\mathbb{R}}^{C\times M}$ is the concatenated hidden state across all $M$ input dimensions, ${\mathbf{C}}_{t}\in{\mathbb{R}}^{1\times C}$, ${\mathbf{D}}\in{\mathbb{R}}^{1\times M}$, and ${\mathbf{u}}_{t}\in{\mathbb{R}}^{1\times M}$. As shown in Fig. 2, the Mamba block used in our work applies a 1-D convolution on the input tokens and then projects the representations to obtain input-dependent SSM parameters [9]. Multiple Mamba blocks are stacked homogeneously. With its selective mechanism [16], Mamba demonstrates strong performance in complex MCL tasks, outperforming other attention-free models and showing competitive results compared to Transformers with key-value caching.
3.2.1 MCL with Mamba as a Continual Learner
We meta-learn Mamba model $f_{\theta}()$ as a continual learner using multiple CL streaming epidodes, where the models learning and prediction as SP relying the retained hidden state within each episode. Each CL episode contains a training data stream ${\mathcal{D}}^{\text{train}}$ and a testing set ${\mathcal{D}}^{\text{test}}$ from the same task distribution, denoted as $P_{({\mathcal{X}},{\mathcal{Y}})}$ with $({\mathcal{D}}^{\text{train}},{\mathcal{D}}^{\text{test}})\sim P_{({\mathcal{X}},{\mathcal{Y}})}$. For instance, in ICL, all testing classes have been seen in the preceding data stream. The objective of MambaCL is to meta-learn the Mamba model’s parameter, i.e., $\theta$, to perform prediction $\hat{y}^{\text{test}}=f_{\theta}(({\mathcal{D}}^{\text{train}},{\mathbf{x}}^{\text{test}}))$ for any $({\mathbf{x}}^{\text{test}}_{i},{\mathbf{y}}^{\text{test}}_{i})\in{\mathcal{D}}^{\text{test}}$. The CL task can be framed as the next token prediction problem in sequence: $({\mathbf{x}}^{\text{train}}_{1},{\mathbf{y}}^{\text{train}}_{1},...,{\mathbf{x}}^{\text{train}}_{T},{\mathbf{y}}^{\text{train}}_{T},{\mathbf{x}}^{\text{test}}_{k})\rightarrow y^{\text{test}}_{k}$. Meta-learning a Mamba continual learner involves optimizing this SP task across multiple sampled CL episodes:
$\min_{\theta}\mathbb{E}_{({\mathcal{D}}^{\text{train}},{\mathcal{D}}^{\text{test}})\sim P_{({\mathcal{X}},{\mathcal{Y}})}}\sum\nolimits_{({\mathbf{x}}^{\text{test}},y^{\text{test}})\in{\mathcal{D}}^{\text{test}}}\ell(f_{\theta}(({\mathcal{D}}^{\text{train}},{\mathbf{x}}^{\text{test}})),y^{\text{test}}),$ (5)
where $\ell(\cdot,\cdot)$ denotes the proper loss function for different tasks, e.g., classification or regression.
Figure 2: The Mamba block in MambaCL.
On the data stream, the meta-learned Mamba $f_{\theta}()$ recognizes the association relationship between ${\mathbf{x}}$ and $y$ through the sequence, recurrently updating the hidden state ${\mathbf{H}}_{t}$ for prediction, as shown in Fig. 1. This efficient online CL process selects and compresses the knowledge in the data stream in a time-variant and content-aware selective manner. To further validate the extension ability of Mamba in MCL, we also explore the potential of incorporating mixture-of-experts (MoE) into the Mamba model [13, 47] for learning and mixing multiple learners (See Sec. 4.3).
Target token embeddings. The value of the target $y$ is essentially a symbol with consistent indication meaning for ${\mathbf{x}}$ within each episode, which does not take any global meaning across the episode. The model is thus trained to handle arbitrary CL episodes with the ability to generalize to different domains. Instead of pre-defining a small and fixed feasible set of candidate targets e.g., classes, and a restricted prediction head, we conduct token embeddings for targets based on a universal and large vocabulary [32], inspired by the tokenization in LMs [54, 12]. For each episode, a subset of unique codes is randomly picked from the vocabulary to indicate different classes; in inference, the sequence model produces the probability of the next step for all possible tokens in the vocabulary. Beyond conducting experiments of meta-training and meta-testing with the same number of classes [32], we conduct generalization analyses in Sec. 4.2.
3.2.2 Selectivity Regularization for Meta Training
Meta-learning a continual learner to associate inputs and targets from a data stream is challenging for both attention-free models and Transformers [32]. Meta-training can be slow and challenging to converge due to long-range dependencies in supervision on the stream. We thus provide additional guidance by strengthening associations (known in the meta-training dataset) between query tokens (i.e., testing inputs) and their correlated preceding tokens. While straightforward for Transformers with explicit intra-sequence attention [32], this process is non-trivial for SSM-based Mamba.
During training, for an input ${\mathbf{x}}$ (corresponding to a pair $({\mathbf{x}},y)$) in the stream at the step $12t+$ 1 after $2t$ tokens of $t$ samples, its association relationship with preceding input tokens can be represented as ${\mathbf{p}}_{2t+1}=[\mathbbm{1}_{y_{2t+1}}(y_{1}),\mathbbm{1}_{y_{2t+1}}(y_{1}),...,\mathbbm{1}_{y_{2t+1}}(y_{t}),\mathbbm{1}_{y_{2t+1}}(y_{t})]\in\{0,1\}^{2t}$, where $\mathbbm{1}_{y}(y^{\prime})$ is an indicator function with $\mathbbm{1}_{y}(y^{\prime})=1,\text{if}~{}y=y^{\prime},~{}\text{and}~{}\mathbbm{1}_{y}(y^{\prime})=0,\text{if}~{}y\neq y^{\prime}$. We hope the meta-learned learner can also identify and use this pattern in CL (i.e., meta-testing).
Transformers store key-value pairs for all samples as state. At each step, explicit attention retrieves learned information by querying all stored keys. As shown in Eq. (1), for the token at step $12t+$ 1, the attention weights/patterns to previous-step tokens can be denoted as ${\mathbf{q}}_{2t+1}^{\text{Trans}}=[{\mathbf{Q}}_{2t+1}{\mathbf{K}}_{j}^{\top}]_{j=1}^{2t}\in{\mathbb{R}}_{\geq 0}^{2t}$. Note that we omit normalization terms in attention weights to simplify the presentation. With explicit attention, meta-learning guidance can be directly applied for Transformers to encourage similarity between ${\mathbf{q}}_{2t+1}^{\text{Trans}}$ and ${\mathbf{p}}_{2t+1}$.
Mamba and other attention-free methods (e.g., Linear Transformer) compress knowledge in a hidden state at each step, as shown in Eq. (2) and (3). Mamba, in particular, applies input-dependent selection and gating at each step. Although Mamba does not generate explicit attention weights, we formulate a regularization for its selectivity by bridging the selective SSM (Eqs.(3) and (4)) with linear attention and softmax attention in Transformers.
As shown in Eq. (2), Linear Transformer updates the state ${\mathbf{S}}$ (and the normalization term ${\mathbf{G}}$) using kernel-based ${\mathbf{K}}$ and ${\mathbf{V}}$, and performs prediction based on ${\mathbf{Q}}$. Considering that the ${\mathbf{K}}$, ${\mathbf{V}}$, and ${\mathbf{Q}}$ in linear attention share the same meaning as in the softmax attention, we still can obtain ${\mathbf{q}}_{2t+1}^{\text{LNTrans}}=[{\mathbf{Q}}_{2t+1}{\mathbf{K}}_{j}^{\top}]_{j=1}^{2t}$ by storing the ${\mathbf{K}}_{j}$ of intermediate tokens only during training for regularization. By examining the duality relationship between the SSM in Eq. (3) and the formulation of Linear Transformer in Eq. (2[9], we can identify the connections between the selective parameters, i.e., ${\mathbf{C}}_{t}$ and ${\mathbf{B}}_{t}$, in SSM and query-key embeddings, i.e., ${\mathbf{Q}}_{t}$ and ${\mathbf{K}}_{t}$, in linear attention. Relying on the linear attention as the bridge, we can obtain the associative indicators of Mamba as ${\mathbf{q}}_{2t+1}^{\text{Mamba}}=[{\mathbf{C}}_{2t+1}{\mathbf{B}}_{j}^{\top}]_{j=1}^{2t}$. To regularize the models’ selection behavior in meta-training, for a query sample $({\mathbf{x}},y)$ in a sequence, we apply a selectivity regularization: $\ell_{\text{slct}}(({\mathbf{x}},y))=\text{KL}({\mathbf{p}}_{\text{idx}(({\mathbf{x}},y))},{\mathbf{q}}^{*}_{\text{idx}(({\mathbf{x}},y))})$, where $\text{idx}()$ indicates the step of the token ${\mathbf{x}}$, * indicates the arbitrary model, and KL divergence is used to minimize the difference between the model’s association pattern and the ground truth. Note that this regularization and maintained intermediate components are not necessary in inference. We apply this regularization to MambaCL and other sequence prediction models (weighted by a scalar $\lambda$) together with the MCL objective in Eq. (5) to enhance the meta-training stability and convergence. Further details are provided in Sec. 6, Sec. C.1, and Sec. C.3 of the Appendix, where we report the $\lambda$ values used in our selectivity regularization loss and present the training loss curves with and without selectivity regularization for different models.
4 Experiments and Analyses
We evaluate different models for MCL and analyze their behavior across various scenarios. Our meta-learning episodes primarily involve classification tasks but also include regression (Sec. 4). Unlike previous studies focused on basic settings, we conduct experiments in both standard MCL setups (Sec. 4) and more diverse configurations (Sec. 4.2).
Experimental setup. For obtaining MCL tasks, given a classification dataset, we firstly divided it into multiple tasks, with each task representing a distinct set of classes. We divide these tasks into two non-overlapping sets, i.e., meta-training and meta-testing. CL episodes in both groups are constructed similarly – each episode randomly, where each episode randomly selects $K$ distinct tasks, with $K$ set to 20 by default. We also explore scenarios with varying $K$ values (Sec. 4.2). By default, each task in both training and testing sequences contains five samples (5-shot). We also explore varying shot numbers to further evaluate adaptability and learning efficiency (Sec. 4.2). For implementation details, please refer to the Appendix Sec. B.
Datasets. We conduct experiments on various datasets, including: general image classification tasks included Cifar-100 [30], ImageNet-1K [53], ImageNet-R [53], MS-Celeb-1M (Celeb) [19], CASIA Chinese handwriting (Casia) [36], and Omniglot [31]; fine-grained recognition tasks involved CUB-200 [62], Stanford Dogs [10], Stanford Cars [29], and FGVC-Aircraft (Aircraft) [40]; the large domain shift tasks featured [46]; and regression tasks consisted of sine wave reconstruction (sine), image rotation prediction (rotation), and image completion (completion). Additional details are provided in the Appendix Sec. A.
Table 1: Classification accuracy (%) across 20-task 5-shot MCL, training from scratch on general image classification tasks. Best and second best rsults are highlighted in red and blue, respectively.
Method Cifar-100 Omniglot Casia Celeb
Meta- Meta- Meta- Meta- Meta- Meta- Meta- Meta-
Train Test Train Test Train Test Train Test
OML $99.4^{\pm 0.1}$ $10.1^{\pm 0.4}$ $99.9^{\pm 0.0}$ $75.2^{\pm 2.2}$ $97.2^{\pm 0.1}$ $96.8^{\pm 0.1}$ $58.2^{\pm 0.3}$ $57.5^{\pm 0.2}$
Transformer $100.0^{\pm 0.0}$ $17.2^{\pm 0.8}$ $100.0^{\pm 0.0}$ $86.3^{\pm 0.6}$ $99.7^{\pm 0.0}$ $99.6^{\pm 0.0}$ $70.9^{\pm 0.2}$ $70.0^{\pm 0.2}$
Linear TF $99.9^{\pm 0.1}$ $16.6^{\pm 0.5}$ $100.0^{\pm 0.0}$ $64.0^{\pm 1.4}$ $99.6^{\pm 0.0}$ $99.3^{\pm 0.0}$ $68.9^{\pm 0.3}$ ${67.6}^{\pm 0.3}$
Performer $100.0^{\pm 0.0}$ $17.1^{\pm 0.3}$ $99.9^{\pm 0.1}$ $62.9^{\pm 4.6}$ $99.5^{\pm 0.0}$ $99.3^{\pm 0.0}$ $67.5^{\pm 0.5}$ $66.3^{\pm 0.2}$
Mamba $99.9^{\pm 0.1}$ $18.3^{\pm 0.4}$ $100.0^{\pm 0.0}$ $87.7^{\pm 0.5}$ $99.8^{\pm 0.1}$ $99.5^{\pm 0.1}$ $69.4^{\pm 0.2}$ $68.1^{\pm 0.1}$
Table 2: Classification accuracy (%) across 20-task 5-shot MCL, training from the pre-trained models on general image classification tasks.
Method Cifar-100 ImageNet-1K ImageNet-R Celeb Casia Omniglot
OML $64.4^{\pm 0.4}$ $90.5^{\pm 0.3}$ $67.5^{\pm 0.3}$ $72.8^{\pm 0.1}$ $81.5^{\pm 0.5}$ $90.4^{\pm 0.2}$
Transformer $62.7^{\pm 0.7}$ $93.5^{\pm 0.1}$ $63.6^{\pm 0.2}$ $78.4^{\pm 0.1}$ $93.8^{\pm 0.2}$ $94.4^{\pm 0.2}$
Linear TF $54.3^{\pm 0.7}$ $89.1^{\pm 0.2}$ $55.7^{\pm 0.3}$ $76.5^{\pm 0.2}$ $90.9^{\pm 0.4}$ $86.5^{\pm 0.5}$
Performer $53.4^{\pm 0.3}$ $90.8^{\pm 0.5}$ $52.8^{\pm 0.9}$ $76.8^{\pm 0.1}$ $93.0^{\pm 0.3}$ $89.3^{\pm 0.3}$
Mamba $67.1^{\pm 0.4}$ $93.6^{\pm 0.2}$ $69.7^{\pm 0.4}$ $77.0^{\pm 0.1}$ $93.1^{\pm 0.2}$ $95.9^{\pm 0.2}$
4.1 Experimental Results and Analyses
We evaluate several models, including OML [25], Vanilla Transformer [65], Linear Transformer [27], Performer [8], and our proposed MambaCL. OML serves as an SGD-based meta-continual learning baseline with a two-layer MLP atop a meta-learned encoder. While Transformers excel at sequence modeling, their computational inefficiency and broad CL objectives limit their suitability for continual learning. Linear Transformer and Performer address this by approximating softmax attention via kernel methods, achieving linear complexity with fixed hidden state size. All transformer variants use 4 layers with 512 hidden dimensions. Mamba, an attention-free model optimized for long sequences, also adopts this architecture in our MambaCL, but with significantly fewer parameters.
General image classification tasks. Table 1 and 2 provide comparative analyses of the performance of various architectures across a range of general image classification tasks, initiating training from scratch and utilizing image representations derived from a pre-trained, respectively. In Table 1, within the CIFAR-100 datasets, all methods suffer from substantial meta-overfitting, as evidenced by the large gap between meta-training and meta-testing scores. This may be attributed to the lower task diversity. In Table 2, continual learners based on pre-trained models show consistent trends on CIFAR-100 and ImageNet-R, with Mamba outperforming other methods, highlighting its robustness against overfitting. On larger datasets like ImageNet-1K, Casia, and Celeb, Mamba matches or exceeds transformer performance. We default to using pre-trained image representations for our experiments to maintain generality. To better understand its behavior, we visualize the associative weights of our meta-trained Mamba Learner in Fig. 3. For further details on the visualization analysis of Transformers and Mamba, please refer to the Appendix Sec. D.
Table 3: Classification accuracy (%) across 20-task 5-shot MCL on fine-grained recognition tasks.
Method CUB-200 Dogs Cars Aircraft
OML $78.7^{\pm 0.6}$ $72.4^{\pm 0.5}$ $83.6^{\pm 0.7}$ $49.5^{\pm 0.2}$
Transformer $81.4^{\pm 0.4}$ $77.5^{\pm 0.6}$ $87.0^{\pm 0.3}$ $53.9^{\pm 0.7}$
Linear TF $69.7^{\pm 0.7}$ $69.7^{\pm 0.7}$ $76.6^{\pm 0.8}$ $49.0^{\pm 0.7}$
Performer $69.2^{\pm 0.8}$ $69.4^{\pm 0.4}$ $73.9^{\pm 0.8}$ $48.6^{\pm 0.6}$
Mamba $83.0^{\pm 0.4}$ $79.2^{\pm 0.5}$ $88.3^{\pm 0.4}$ $55.3^{\pm 0.6}$
Fine-grained recognition tasks. Table 3 presents a performance comparison of different architectures on fine-grained recognition datasets. In fine-grained datasets characterized by subtle inter-class differences (e.g., the CUB-200 dataset, which contains 200 bird subcategories), models need to integrate global information throughout the training episode to identify these nuances effectively. Mamba outperforms other models across these datasets, potentially attributable to its robustness to capture subtle inter-class distinctions.
Table 4: Classification accuracy (%) and regression errors across 100-task 5-shot MCL.
Method Casia Celeb Sine Rotation Completion
OML $93.2^{\pm 0.9}$ $45.5^{\pm 0.2}$ $0.0498^{\pm 0.0004}$ $0.524^{\pm 0.087}$ $0.1087^{\pm 0.0001}$
Transformer $99.0^{\pm 0.0}$ ${60.5}^{\pm 0.1}$ ${0.0031}^{\pm 0.0002}$ ${0.031}^{\pm 0.001}$ ${0.0989}^{\pm 0.0001}$
Linear TF $97.7^{\pm 0.1}$ $54.7^{\pm 0.1}$ ${0.0139}^{\pm 0.0003}$ ${0.047}^{\pm 0.002}$ ${0.1084}^{\pm 0.0001}$
Mamba $99.1^{\pm 0.1}$ $59.9^{\pm 0.1}$ $0.0054^{\pm 0.0001}$ $0.025^{\pm 0.001}$ $0.0895^{\pm 0.0001}$
Training on Longer Episodes. We conducted experiments to meta-train the models on longer episodes across both classification and regression tasks. Table 4 demonstrates that Mamba continues to perform comparably to Transformer, and significantly outperforms SGD-based approaches (OML).
Figure 3: Final-layer associations in a 20-task 5-shot meta-testing episode of MambaCL (meta-trained on 20-task 5-shot MCL). All three visualizations share the same training episode (shots $0^{th}\!\!-\!99^{th}$), with test queries at the $100^{th}$ shot from the $1^{st}$, $9^{th}$, and $18^{th}$ tasks, aligning with the $5^{th}\!\!-\!9^{th}$, $45^{th}\!\!-\!49^{th}$, and $90^{th}\!\!-\!94^{th}$ training shots. The red box highlights each query and its corresponding examples. See Appendix Sec. D for additional visualizations comparing Transformer attention and Mamba’s associative selectivity.
4.2 Generalization Analyses
We hope a meta-learned learner has the ability to be generalized to unseen scenarios. To assess this, we conduct generalization analyses for Transformer models and Mamba in conditions involving longer untrained sequence lengths, significant domain shifts, and sensitivity to noisy inputs during meta-testing. Additionally, to understand the behavior of these models, we visualize the attention weights of Transformers and the associative weights of Mamba, illustrating their attention mechanisms and selectivity patterns in the Appendix Sec. D.
Generalization to varying sequence length. To address continual learning episodes of indefinite length effectively, the learning algorithm should demonstrate the capability to generalize beyond the sequence lengths encountered during meta-training. We conducted length generalization experiments on ImageNet-1K, training vanilla and linear Transformers, and Mamba on 20-task 5-shot MCL, each with a vocabulary of 200 tokens. The length of a continual learning episode is calculated as $2\times\text{tasks}\times\text{shots}+1$.
(a) Meta-testing on varying numbers of tasks. Fig. 5 illustrates the performance of the three models, meta-trained on a 20-task 5-shot setup and evaluated during meta-testing across varying numbers of tasks while keeping a constant shot number of 5. Both the vanilla Transformer and Linear Transformer suffer significant performance degradation during meta-testing at untrained episode lengths, even in configurations such as the 10-task 5-shot setup. Mamba demonstrates superior performance in the 10-task setup compared to the meta-trained 20-task setup, and its performance degradation is relatively mild compared to transformers as the number of tasks increases.
(b) Meta-testing on varying numbers of shots. In Fig. 5, we assess the performance of three models meta-trained in a 20-task 5-shot setup and evaluated during meta-testing across varying shots numbers while keeping a constant task number of 20. Both the vanilla Transformer and linear Transformer suffer significant performance degradation, likely due to overfitting the 20-task 5-shot pattern. In contrast, Mamba exhibits only about a 10% performance degradation when the number of shots in meta-testing increases to 50, which is ten times the length of the meta-training episodes. Fig. 5 and 5 highlight Mamba’s robustness in adapting to longer sequence lengths (length generalization).
(a) Task
(b) Shot
(c) Noise
Figure 4: Generalization analysis on ImageNet-1K, with meta-training on 20 tasks (5-shot), meta-testing on: (a) varying tasks (5-shot), (b) varying shots (20-task), and (c) varying input noise levels (20-task 5-shot).
(a) $\lambda$
(b) SSM state size
Figure 5: Ablation studies: (a) varying $\lambda$ in training loss; (b) varying SSM state size.
Table 5: Classification accuracy (%) across 20-task 5-shot MCL on DomainNet dataset. (inf,pnt,qdr,rel,skt→clp denotes meta-testing on the Clipart domain, and the remaining domains used for meta-training. clp: clipart, inf: infograph, pnt: painting, qdr: quickdraw, rel: real, skt: sketch.)
Method inf,pnt,qdr, clp,pnt,qdr, clp,inf,qdr, clp,inf,pnt, clp,inf,pnt, clp,inf,pnt, Avg
rel,skt→clp rel,skt→inf rel,skt→pnt rel,skt→qdr qdr,skt→rel qdr,rel→skt
Transformer $91.8^{\pm 0.1}$ $69.4^{\pm 0.1}$ $82.6^{\pm 0.2}$ $50.2^{\pm 0.6}$ $93.8^{\pm 0.1}$ $85.9^{\pm 0.3}$ $79.0^{\pm 0.3}$
Linear TF $91.0^{\pm 0.0}$ $66.2^{\pm 0.8}$ $80.6^{\pm 0.8}$ $30.7^{\pm 1.4}$ $92.9^{\pm 0.1}$ $85.5^{\pm 0.1}$ $74.5^{\pm 0.5}$
Performer $91.3^{\pm 0.2}$ $66.4^{\pm 0.6}$ $81.3^{\pm 0.2}$ $39.4^{\pm 1.7}$ $92.8^{\pm 0.1}$ $84.8^{\pm 0.5}$ $76.0^{\pm 0.6}$
Mamba $91.7^{\pm 0.2}$ $70.2^{\pm 0.2}$ $81.8^{\pm 0.2}$ $55.6^{\pm 0.8}$ $93.0^{\pm 0.1}$ $87.2^{\pm 0.3}$ $79.9^{\pm 0.3}$
Results and analyses on larger domain shift. We investigate a larger domain shift scenario using the DomainNet dataset, which includes six distinct domains, to further assess model generalization to unseen input distributions, a setting that mirrors real-world conditions. One domain is held out for meta-testing, while the others are used for meta-training. The experimental results are detailed in Table 5. Overall, these models demonstrate the capability to handle large domain shift scenarios effectively. Mamba performs on par with or surpasses transformers across diverse target domains, benefiting from potentially enhanced generalization capabilities associated with its smaller size and reduced susceptibility to overfitting. Vanilla Transformers perform well with the targets are real images or paintings. Mamba excels in the quickdraw domain, which presents greater variances from other domains. This performance advantage may be ascribed to Mamba’s robustness in handling inputs that deviate significantly from the training distribution.
Sensitivity to the noisy inputs. To assess the sensitivity of different models to noisy inputs, we conduct experiments on meta-trained 20-task 5-shot MCL models using the ImageNet-1K dataset. During each meta-testing episode, we apply noise to the input embeddings ${\mathbf{x}}_{i}$ of five randomly selected samples, adhering to Gaussian distributions with a mean ($\mu$) of 00 and a standard deviation ($\sigma$) varying from 00 to $10$. As illustrated in Fig. 5, the vanilla transformer and linear transformer suffer significant performance degradation. In contrast, Mamba demonstrates robust performance, effectively handling inputs with increased noise levels/strength.
4.3 Ablation Studies
Table 6: Different Mamba architectures on 20-task 5-shot MCL.
Method Cifar-100 ImageNet-1K
Transformer $62.7^{\pm 0.7}$ $93.5^{\pm 0.1}$
Linear TF $54.3^{\pm 0.7}$ $89.1^{\pm 0.2}$
Mamba-1 $59.7^{\pm 0.5}$ $90.1^{\pm 0.3}$
MambaFormer $62.4^{\pm 0.6}$ $92.7^{\pm 0.1}$
Mamba-2 $67.1^{\pm 0.4}$ $93.6^{\pm 0.2}$
Mamba+MoE $68.9^{\pm 0.2}$ $94.0^{\pm 0.2}$
Hyper-parameter of selectivity regularization loss. We conducted ablation studies to assess the impact of the selectivity regularization weight $\lambda$ on our Mamba model’s performance, varying it across ${0.1,0.2,0.5,1.0,2.0}$ (Fig. 5). The results indicate that performance remains largely stable across different $\lambda$ values, suggesting Mamba’s robustness to this hyper-parameter. Based on these findings, we set $\lambda=0.5$ for all reported experiments.
SSM state size. In Fig. 5, we assess the impact of varying SSM state size on the performance of our methods. We conducted experiments on ImageNet-1K and Cifar-100, training the MambaCL with state sizes of 16, 32, 64, 128, and 256. The results demonstrate a consistent improvement in performance as the state size increases. To optimally balance performance with computational efficiency, we selected a state size of 128 for our experiments.
Different architectures. Table 6 presents an ablation study comparing Mamba-1, MambaFormer [45], and Mamba-2 within our MambaCL framework. MambaFormer is a hybrid model that integrates vanilla attention into Mamba and replaces Transformer positional encoding with a Mamba block. The results show MambaFormer performs comparably to Transformers, while Mamba-2 achieves the best performance on the Cifar-100 dataset.
Mamba+MoE. In Table 6, we present experiments in which Mamba is augmented with mixture-of-experts (MoE), incorporating twelve 2-layer MLP expert networks with a dense-MoE router following each Mamba Block, resulting in improved performance. Additionally, we include performance metrics for vanilla and linear transformers as references.
Table 7: Computational cost on 20-task 5-shot MCL.
[…]
TF $9.2$M $325$ep/s
Mamba $5.4$M $858$ep/s
Computational cost. In Table 7, we detail various aspects of computational cost using our implementation in PyTorch, executed on an NVIDIA 4090 and an INTEL I9-14900K. We specifically report the costs associated with meta-testing at a batch size of 1. Mamba achieves comparable or superior performance to the vanilla Transformer, while being more efficient in parameters and speed.
5 Conclusion
In this paper, we try to answer the question – Can attention-free Mamba perform well on MCL? We formulate SSMs and Mamba as a sequence prediction-based continual learner and meta-learn it on CL episodes. A selectivity regularization is introduced for meta-continual learning the models. Comprehensive experiments show that Mamba performs well across diverse MCL scenarios, significantly outperforming other attention-free methods and matching or exceeding Transformers’ performance with fewer parameters and computations. In challenging scenarios with global structures, domain shifts, and long sequences, Mamba demonstrates obvious reliability, generalization, and robustness.
Limitations and future work. This study can be extended to larger-scale datasets and offline CL settings. Future work also includes exploring other types efficient models, such as efficient Transformers with compressed key-value caches and efficient CL for large-scale foundation models. Beyond the current MCL framework, we aim to expand its applicability to a broader range of scenarios and models.
Appendix A Datasets
A.1 General Image Classification Tasks
Cifar-100 [30] dataset consists of 60,000 images across 100 classes, each with 600 images. We select 60 classes at random for meta-training and use the remaining 40 for meta-testing.
ImageNet-1K [53] dataset comprises over one million labeled images distributed across 1,000 categories. We select 600 classes at random for meta-training and use the remaining 400 for meta-testing.
ImageNet-R [53] extends 200 ImageNet classes with a compilation of 30,000 images tailored for robustness research.
Celeb [19] is a large-scale facial image collection featuring approximately 10 million images of 100,000 celebrities. We randomly allocated 1,000 classes for meta-testing and assigned the remaining classes to meta-training.
Casia Chinese handwriting [36] dataset encompasses a total of 7,356 character classes with 3.9 million images. We randomly selected 1,000 classes for the meta-testing and allocated the remaining classes for meta-training.
Omniglot [31] is a collection of 1,632 handwritten characters from 50 different alphabets. The meta-training set comprises 963 classes, while the meta-testing set includes 660 classes, with each class containing 20 images.
A.2 Fine-grained Recognition Tasks
CUB-200-2011 [62] is a widely used fine-grained visual categorization dataset, comprising 11,788 images across 200 bird subcategories. We randomly selected 80 classes for the meta-testing and allocated the remaining classes for meta-training.
Stanford Dogs [10] dataset comprises 20,580 images spanning 120 global dog breeds, divided into 12,000 training images and 8,580 testing images. We select 48 classes at random for meta-testing and use the remaining 72 for meta-training.
Stanford Cars [29] comprises 16,185 images across 196 car classes, primarily captured from the rear perspective. We select 80 classes at random for meta-testing and use the remaining 40 for meta-training.
FGVC-Aircraft [40] dataset comprises 10,200 images across 102 aircraft model variants, each represented by 100 images, primarily consisting of airplanes. We randomly selected 40 classes for the meta-testing and allocated the remaining classes for meta-training.
A.3 Large Domain Shift Tasks
DomainNet [46] dataset is a benchmark for domain adaptation, encompassing common objects organized into 345 classes across six domains: clipart, real, sketch, infograph, painting, and quickdraw. We evaluate model adaptability to out-of-domain data by using one domain for meta-testing and the remaining domains for meta-training.
A.3.1 Regression Tasks
Sine Wave Reconstruction (Sine) The sine wave $\omega(\tau)=A\sin(2\pi\nu\tau+\psi)$ is defined by its amplitude $A$, frequency $\nu$, and phase $\psi$. We denote the target values $y$ as evaluations of the sine wave at 50 predefined points: $y=[\omega(\tau_{1}),\dots,\omega(\tau_{50})]$. In each task, the frequency and phase remain constant, but the amplitude is allowed to vary. To corrupt $y$ into $x$, we introduce a phase shift and Gaussian noise, where the phase shift is randomly selected for each task. The mean squared error between $y$ and the model’s prediction $\hat{y}$ is reported as the evaluation criterion.
Image Rotation Prediction (Rotation) The model is provided with an image rotated by an angle $\psi\in[0,2\pi)$, and its task is to predict the rotation angle $\hat{\psi}$. We use $1-\cos(\hat{\psi}-\psi)$ as the evaluation metric, where a perfect prediction would result in a score of 0, while random guessing would yield an average score of 1.0. The Casia dataset is employed, with each class being treated as an individual task, maintaining the same meta-split configuration.
Mamba Transformer Linear TF Performer
Batch size 16
Max Train Step 50000
Optimizer Adam
Learning Rate $1\times 10^{-4}$
Learning Rate Decay Step
Learning Rate Decay Step 10000
Learning Rate Decay Rate 0.5
Regularization $\lambda$ 0.5
Hidden Dimension 512
Layer 4
State Size 128 - - -
Delta Convolution 4 - - -
Attention - Softmax Elu Favor
Table 8: Model Configurations
(a) Mamba w/ $\ell_{\text{slct}}$
(b) Transformer w/ $\ell_{\text{slct}}$
(c) Linear TF w/ $\ell_{\text{slct}}$
(d) Performer w/ $\ell_{\text{slct}}$
(e) Mamba w/o $\ell_{\text{slct}}$
(f) Transformer w/o $\ell_{\text{slct}}$
(g) Linear TF w/o $\ell_{\text{slct}}$
(h) Performer w/o $\ell_{\text{slct}}$
Figure 6: Training loss curves for (a, e) Mamba, (b, f) Transformer, (c, g) Linear Transformer, and (d, h) Performer, under the same type of representation and experimental settings, with and without selectivity regularization ($\ell_{\text{slct}}$) during meta-training on 20-task, 5-shot MCL on Cifar-100.
Image Completion (Completion) In this task, the model is tasked with filling in the missing parts of an image given the visible sections. Using the Casia dataset, we modify the input $x$ to consist of the top half of the image, while the target $y$ is the bottom half. The performance is evaluated by computing the mean squared error between the predicted and true pixel values. We report the MSE between $y$ and the model’s prediction $\hat{y}$ as the evaluation criterion.
Appendix B Implementation Details
We conduct our main experiments on a single NVIDIA A100 GPU, repeating each experiment five times and reporting the mean and standard deviation. Results are recorded upon convergence on the meta-training set. We use a batch size of 16 and the Adam optimizer, with an initial learning rate of $1\times 10^{-4}$ that decays by $0.5$ every $10,000$ steps. All models follow a consistent setup to ensure fair comparisons, with details provided in the Appendix. For experiments involving training from scratch, we adopt settings from [32] for fairness. For experiments with pre-trained networks, we employ OpenAI/CLIP-ViT-B16 [48, 24] as the image encoder, keeping its parameters frozen and adding a trainable linear projector. Table 8 presents the configurations of the models employed in our experiments.
(a) Mamba
(b) Transformer
(c) Linear TF
(d) Performer
Figure 7: Ablation studies on regularization strength $\lambda$ (${0.1,0.2,0.5,1.0,2.0}$) during meta-testing of 20-task, 5-shot models (meta-trained on 20-task, 5-shot) for (a) Mamba, (b) Transformer, (c) Linear Transformer, and (d) Performer.
Appendix C Additional Experiments
C.1 Effects of Selectivity Regularization and Meta-Training Loss Curves
Due to the complexity of the MCL task, the regularization technique plays a crucial role in stabilizing and improving the training process. Fig. 6 showing the initial training phases (2500 steps) for different models with and without selectivity regularization. The losses are 3–5 times higher compared to the models with regularization applied and successfully converging. Beyond 2500 steps, the losses oscillate and no longer decrease. The results indicate that models without our regularization struggle to converge and exhibit significant oscillations during training, highlighting the effectiveness of the regularization.
C.2 Ablation Studies on Regularization Strength
Fig. 5 in the main paper, we conducted an ablation study to assess the influence of regularization strengths on our Mamba’s efficacy. Fig. 7 illustrates more ablation studies assessing the impact of regularization strengths $\lambda$ by setting is as $\{0.1,0.2,0.5,1.0,2.0\}$, across multiple models on both ImageNet-1K and Cifar-100 datasets. The results demonstrate that all models exhibit stability within a wide and appropriate range of $\lambda$, providing evidence of consistent patterns. In our experiments, without losing generality, all models employed a regularization strength of 0.5 by default.
C.3 Ablation Studies on Learning Rates
Fig. 8 illustrates ablation studies assessing the impact of varying initial learning rates $\{5\times 10^{-5},1\times 10^{-4},2\times 10^{-4},5\times 10^{-4}\}$, across multiple models on both ImageNet-1K and Cifar-100 datasets. The results indicate that within a reasonable range, the learning rate does not significantly affect model performance. In our experiments, without losing generality, we set the initial learning rate to $1\times 10^{-4}$, with decays of 0.5 every 10,000 steps.
(a) Mamba
(b) Transformer
(c) Linear TF
(d) Performer
Figure 8: Ablation studies on learning rates ($\{5\times 10^{-5},1\times 10^{-4},2\times 10^{-4},5\times 10^{-4}\}$) during meta-testing of 20-task, 5-shot models (meta-trained on 20-task, 5-shot) for (a) Mamba, (b) Transformer, (c) Linear Transformer, and (d) Performer.
C.4 Additional Generalization Analyses
Without the regularization, models struggle to converge and exhibit significant oscillations during training, as shown in Fig. 6. In Sec. 4.2 and Fig. 5 of the main paper, we conducted generalization analyses of various models by conducting meta-testing on the episodes different from the meta-training settings. Specifically, we apply the models meta-trained with 20-task-5-shot episodes on the meta-testing episodes with varying numbers of tasks or shots or the episodes contaminated by noise. The results show that Mamba shows better generalization ability to unseen scenarios and Transformer shows more meta-overfitting issues. To validate that the results are not relevant to the regularization, we evaluated various models with a small regularization strength ($\lambda=0.1$) to assess the impact of regularization on this generalization experiment and the meta-overfitting issue. The results indicate that regularization strengths of 0.1 (9) and 0.5 (5 in the main paper) lead to similar phenomena across different models.
(e) Task
(f) Shot
(g) Noise
Figure 9: Generalization Analysis on ImageNet-1K with regularization strength $\lambda=0.1$: (a) meta-trained on 20-task 5-shot MCL, meta-testing on varying number of tasks (5-shot); (b) meta-trained on 20-task 5-shot MCL, meta-testing on varying number of shots (20-task); (c) meta-trained on 20-task 5-shot MCL, meta-testing on 20-task 5-shot with varying inputs noise intensity level
Appendix D Visualization of Attention and Selectivity Pattern
Given meta-learned sequence models as the continual learner, the models process the samples in sequence in the meta-test CL process. To analyze the behaviors of these models, we visualize the attention weights of Transformers and the associative weights of Mamba (as discussed in 3.2.2 in the main paper) to demonstrate their attention and selectivity patterns, respectively. In a meta-testing episode, given a trained model and a sequence of samples, the prediction for a given ${\mathbf{x}}^{\text{test}}$ is produced based on the attention or implicit association of seen samples in the sequence. Visualizing the attention and selectivity patterns can empirically show how the models make predictions. For the standard benchmarking case, Fig. 10 shows that both Transformer and Mamba can effectively associate seen samples with query inputs, leading to the results as shown in Table 2 in the main paper.
Specifically, we use this visualization to analyze how different models perform in the generalization studies (discussed in 4.2 of the main paper), i.e., generalizing to meta-testing cases that are different from meta-training cases.
D.1 Visualization Analyses for Generalization to Different Stream Length
The experiments shown in 5 and 5 in the main paper validate the generalization ability of models by meta-testing on CL episodes/sequences that differ from those seen during meta-training. Specifically, the models are meta-trained on 20-task, 5-shot MCL episodes and meta-tested on episodes with task and shot numbers exceeding those in meta-training. Transformers generally converge more easily during meta-training compared to Mamba, due to their strong fitting ability. However, this advantage may also lead to meta-overfitting.
To analyze how different models perform on these sequences, we visualize the final layer attention weights of Transformer and the corresponding selective scores (associative indicators) of Mamba, between various test shots (queries) and a single MCL train episode (prompt) of both Mamba and Transformer. Note that Mamba does not have explicit attention weights, we compute the scores relying on the connection between Mamba and Transformers described in 3.2.2 in the main paper. Specifically, we computed the parameters ${\mathbf{C}}_{\text{test}}$ and ${\mathbf{B}}$ (${\mathbf{C}}_{\text{test}}{\mathbf{B}}^{\top}$) within its SSMs to compare its behavior with the attention matrix (${\mathbf{Q}}_{\text{test}}{\mathbf{K}}^{\top}$) of Transformers, where ${\mathbf{C}}_{\text{test}}\in{\mathbb{R}}^{1\times C}$ and ${\mathbf{Q}}_{\text{test}}\in{\mathbb{R}}^{1\times C}$ correspond to the row of the test shot. Both models are meta-trained on a 20-task, 5-shot setting using the ImageNet-1K dataset.
For models meta-trained on the 20-task, 5-shot setting, we meta-tested them and visualized their weights on 20-task, 5-shot episodes (Fig. 10), 20-task, 10-shot episodes (Fig. 11), and 40-task, 5-shot episodes (Fig. 12). Specifically, we observed that Transformers tend to either average attention or consistently focus on specific token positions in episodes that deviate from the training length. In contrast, Mamba effectively associates with relevant shots. This suggests that Transformers may learn pattern biases in the sequences (e.g., positional biases unrelated to content), leading to meta-overfitting during these generalization tests.
(a) Mamba
(b) Transformer
Figure 10: 20-task 5-shot in meta-testing: visualization of the final layer associations between various test shots (queries) and a single MCL train episode (prompt) of both (a) Mamba and (b) Transformer during meta-testing on 20-task 5-shot MCL episode (meta-trained on 20-task 5-shot). In meta-testing, the four visualizations share a single MCL training episode (prompt) spanning $0^{th}\!\!-\!99^{th}$ shots, while the test shots (queries at the $100^{th}$ shot) correspond to the $0^{th}$, $1^{st}$, $9^{th}$, and $18^{th}$ tasks ($0^{th}\!\!-\!4^{th}$, $5^{th}\!\!-\!9^{th}$, $45^{th}\!\!-\!49^{th}$, and $90^{th}\!\!-\!94^{th}$ train shots), respectively.
(a) Mamba
(b) Transformer
Figure 11: More shots in meta-testing: visualization of the final layer associations between various test shots (queries) and a single MCL train episode (prompt) of both (a) Mamba and (b) Transformer during meta-testing on 20-task 10-shot MCL episode (meta-trained on 20-task 5-shot). In meta-testing, the four visualizations share a single MCL training episode (prompt) spanning $0^{th}\!\!-\!199^{th}$ shots, while the test shots (queries at the $100^{th}$ shot) correspond to the $0^{th}$, $1^{st}$, $9^{th}$, and $18^{th}$ tasks, respectively.
(a) Mamba
(b) Transformer
Figure 12: More tasks in meta-testing: visualization of the final layer associations between various test shots (queries) and a single MCL train episode (prompt) of both (a) Mamba and (b) Transformer during meta-testing on 40-task 5-shot MCL episode (meta-trained on 20-task 5-shot). In meta-testing, the seven visualizations share a single MCL training episode (prompt) spanning $0^{th}\!\!-\!199^{th}$ shots, while the test shots (queries at the $100^{th}$ shot) correspond to the $0^{th}$, $1^{st}$, $9^{th}$, $18^{th}$, $28^{th}$, $38^{th}$and $39^{th}$ tasks, respectively.
D.2 Visualization Analysis of Generalization to Noise-Contaminated Episodes
(a) Mamba
(b) Transformer
Figure 13: Noise inputs in meta-testing: visualization of the final layer associations between various test shots (queries) and a single MCL train episode (prompt) of both (a) Mamba and (b) Transformer (meta-trained on 20-task 5-shot without noise inputs), during meta-testing on 20-task 5-shot MCL episode with noise inputs (noise strength=1, noise on $8^{th}$, $18^{st}$, $39^{th}$, $61^{th}$, and $75^{th}$ training shots). In meta-testing, the four visualizations share a single MCL training episode (prompt) spanning $0^{th}\!\!-\!99^{th}$ shots, while the test shots (queries at the $100^{th}$ shot) correspond to the $0^{th}$, $1^{st}$, $9^{th}$, and $18^{th}$ tasks ($0^{th}\!\!-\!4^{th}$, $5^{th}\!\!-\!9^{th}$, $45^{th}\!\!-\!49^{th}$, and $90^{th}\!\!-\!94^{th}$ train shots), respectively.
(a) Mamba
(b) Transformer
Figure 14: Noise inputs in meta-testing: visualization of the final layer associations between various test shots (queries) and a single MCL train episode (prompt) of both (a) Mamba and (b) Transformer (meta-trained on 20-task 5-shot without noise inputs), during meta-testing on 20-task 5-shot MCL episode with noise inputs (noise strength=2, noise on $8^{th}$, $18^{st}$, $39^{th}$, $61^{th}$, and $75^{th}$ training shots). In meta-testing, the four visualizations share a single MCL training episode (prompt) spanning $0^{th}\!\!-\!99^{th}$ shots, while the test shots (queries at the $100^{th}$ shot) correspond to the $0^{th}$, $1^{st}$, $9^{th}$, and $18^{th}$ tasks ($0^{th}\!\!-\!4^{th}$, $5^{th}\!\!-\!9^{th}$, $45^{th}\!\!-\!49^{th}$, and $90^{th}\!\!-\!94^{th}$ train shots), respectively.
(a) Mamba
(b) Transformer
Figure 15: Noise inputs in meta-testing: visualization of the final layer associations between various test shots (queries) and a single MCL train episode (prompt) of both (a) Mamba and (b) Transformer (meta-trained on 20-task 5-shot without noise inputs), during meta-testing on 20-task 5-shot MCL episode with noise inputs (noise strength=6, noise on $8^{th}$, $18^{st}$, $39^{th}$, $61^{th}$, and $75^{th}$ training shots). In meta-testing, the four visualizations share a single MCL training episode (prompt) spanning $0^{th}\!\!-\!99^{th}$ shots, while the test shots (queries at the $100^{th}$ shot) correspond to the $0^{th}$, $1^{st}$, $9^{th}$, and $18^{th}$ tasks ($0^{th}\!\!-\!4^{th}$, $5^{th}\!\!-\!9^{th}$, $45^{th}\!\!-\!49^{th}$, and $90^{th}\!\!-\!94^{th}$ train shots), respectively.
In the experiments, the modes are meta-trained on noise-free episodes. The noise is added to randomly selected samples/shots in the meta-testing episodes. The task can also be seen as validating the ability to ignore the irrelevant samples or contaminated outlier samples in the sequences. To directly show how the models work in this scenario, we visualized the final layer attention weights for test shots compared to training shots for both Mamba and Transformer, each meta-trained in a 20-task, 5-shot setting. During meta-testing, these models processed a 20-task, 5-shot episode with five noisy input shots (shot index: 8, 18, 39, 61, 75) at noise strengths of 1 (Fig. 13), 2 (Fig. 14), and 6 (Fig. 15). The results indicate that the Transformer meta-trained on clean episodes tends to produce extreme attention weights (either very high or very low) on noisy or outlier shots, whereas Mamba is less affected. This observation suggests that Transformers’ learned attention mechanisms tend to associate samples based on local and independent representations. In contrast, Mamba performs more effectively by selectively associating relevant information and leveraging its recurrently updated latent state, which accumulates global sequence information.
Appendix E Broader impacts
This work presents both promising and cautionary societal implications. On the positive side, MambaCL enables efficient continual learning with low memory and computation costs, making it well-suited for deployment in resource-constrained or privacy-sensitive environments such as mobile devices, robotics, and personalized healthcare systems. Its robustness to domain shifts and noisy inputs further supports real-world reliability. However, our model has not been trained or evaluated on larger-scale datasets or in real-world industry scenarios, so its behavior and potential impacts in such settings remain uncertain. Care should be taken when applying it to more complex or practical environments.